# Identifying Legal Holdings with LLMs: A Systematic Study of Performance, Scale, and Memorization

### v2

@@ Abstract.
Large Language Models, Natural Language Processing, Legal Analytics, Judicial Reasoning
copyright: acmcopyrightjournalyear: 2025conference: International Conference on Artificial Intelligence and Law 2025; June 16 - 20, 2025; Chicago, IL
@@ 2.1. CaseHOLD Dataset
Figures 1 and 2 present two example questions from this dataset of varying difficulty. The input prompt is shown on the left, and the model must identify the correct choice to complete the parenthetical ¡HOLDING¿ citation. On the right are the five answer choices, with the correct answer choice highlighted in blue, and the models which selected that answer are shown in brackets. In some cases, like 1, key terms like “Coast Guard” and “vessel” hint at the correct answer, while in others like 2, the model must select from multiple legal doctrines, including jurisdictional quuestion and multiple flavors of the “filed rate” doctrine.
@@
[style=mystyle] and suppressing violations of laws of the United States, “officers may at any time go on board of any vessel subject to the jurisdiction, or to the operation of any law, of the United States, address inquiries to those on board, examine the ship’s documents and papers, and examine, inspect, and search the vessel and use all necessary force to compel compliance.” This statute has been construed to permit the Coast Guard to stop an American vessel in order to conduct “a document and safety inspection on the high seas, even in the absence of a warrant or suspicion of wrongdoing,” United States v. Hilton, 619 F.2d 127, 131 (1st Cir.1980), and to conduct a more intrusive search on the basis of reasonable suspicion, see United States v. Wright-Barker, 784 F.2d 161, 176 (3d Cir.1986) (¡HOLDING¿), superseded by statute on other grounds as holding that forfeiture statute is subject to the fourth amendments prohibitions against unreasonable searches and seizures holding that the fourth amendment proscription against unreasonable searches and seizures was applicable to the states under the fourteenth amendment so that evidence seized in violation of the constitution could no longer be used in state courts holding sbm is not a violation of the defendants fourth amendment right to be free from unreasonable searches and seizures holding that a reasonable suspicion requirement for searches and seizures on the high seas survives fourth amendment scrutiny[ALL MODELS] holding that inmates fourth amendment protection from unreasonable strip searches survives hudson
@@ 2.3. Zero-Shot Empirical Results
First, we measure the diversity of the model responses. When two models respond with the same answer to a given question, we code that as a match. The match rate then measures how often two models give the same answer over all questions. In Figure 4, we show the match rate between the answers produced by all seven models. For example, the GPT-4o and NovaMicro models produced the same answer in 70.0% of cases. All model pairs produced the same answer more than 50% of the time.
@@
We also measure the number of unique choices to a given question. Table 1 shows the number of unique choices for each question. In 34% of cases, all models tested chose the same answer - for these questions, the models identified the correct choice more than 90% of the time. In other cases, the models split on the correct answer. Accuracy for questions with two or more different answers was much lower.
@@
As in (Zheng et al., 2021), we compute the macro F1 score which measures both precision (the fraction of correct predictions among all predictions of a given class) and recall (the fraction of actual instances of a class that were correctly identified). The macro F1 score averages the F1 score across all five answer choice classes to account for potential imbalance in how often each appears. Higher values are better, with 1 representing perfect performance. We include the best performing reported results from three papers examining the CaseHOLD dataset (Zheng et al., 2021; Chalkidis et al., 2022; Niklaus et al., 2024).
@@
  • As Figure 5 shows, performance scales with model size. Among the Llama, Nova and GPT-4o model families, performance on this task improves with parameter size. This aligns with the scaling laws seen elsewhere in the literature (for example, (Brown et al., 2020; Kaplan et al., 2020)), and suggests that improvements to general purpose LLMs may also improve performance on specific legal tasks.
  • Several of these models match or surpass the performance of custom-built legal models. For example, GPT-4o performance of 0.744 surpasses the best reported results in (Zheng et al., 2021), 5 of the 7 reported results in (Chalkidis et al., 2022) (ranging from 0.708 to 0.754) and 3 of the 5 reported models presented in (Niklaus et al., 2024) (ranging from 0.717 to 0.770). These models do this without “fine-tuning” or domain adaptation. This means that general purpose LLMs may now take on tasks that historically required expensive and sophisticated fine-tuning.
@@ 3. Have LLMs memorized the test set?
To do so, we employ a two-step prompting scheme. In the first step, we use Prompt 6 to prompt the GPT4o-mini model to anonymize the citation in question, replacing case names and citations with similar but artificial citations. The resulting text is semantically similar to the original, but modifications ensure that the text is novel and that the citations themselves are meaningless. An example of such an original and modified citation prompt is presented in Figure 7. Note that all names and citations have been replaced with plausible alternatives, while the underlying facts and structure is unchanged.

### v3

@@ Abstract.
Large Language Models, Natural Language Processing, Legal Analytics, Judicial Reasoning, Memorization, Model Scaling
copyright: acmcopyrightjournalyear: 2025conference: International Conference on Artificial Intelligence and Law 2025; June 16 - 20, 2025; Chicago, ILccs: Applied computing Lawccs: Computing methodologies Information extractionccs: Computing methodologies Natural language processing
@@ 2.1. CaseHOLD Dataset
Figures 1 and 2 present two questions from this dataset of varying difficulty. The input prompt is on the left, and the model must identify the correct choice to complete the parenthetical ¡HOLDING¿ citation. On the right are the five answer choices, with the correct answer choice highlighted in blue, and the models which selected that answer in brackets. In Figure 1, key terms like “Coast Guard” and “vessel” hint at the correct answer, while in Figure 2, the model must select from multiple legal doctrines, including jurisdictional questions and multiple flavors of the “filed rate” doctrine.
@@
and suppressing violations of laws of the United States, “officers may at any time go on board of any vessel subject to the jurisdiction, or to the operation of any law, of the United States, address inquiries to those on board, examine the ship’s documents and papers, and examine, inspect, and search the vessel and use all necessary force to compel compliance.” This statute has been construed to permit the Coast Guard to stop an American vessel in order to conduct “a document and safety inspection on the high seas, even in the absence of a warrant or suspicion of wrongdoing,” United States v. Hilton, 619 F.2d 127, 131 (1st Cir.1980), and to conduct a more intrusive search on the basis of reasonable suspicion, see United States v. Wright-Barker, 784 F.2d 161, 176 (3d Cir.1986) (¡HOLDING¿), superseded by statute on other grounds as
  • holding that forfeiture statute is subject to the fourth amendments prohibitions against unreasonable searches and seizures
  • holding that the fourth amendment proscription against unreasonable searches and seizures was applicable to the states under the fourteenth amendment so that evidence seized in violation of the constitution could no longer be used in state courts
  • holding sbm is not a violation of the defendants fourth amendment right to be free from unreasonable searches and seizures
  • holding that a reasonable suspicion requirement for searches and seizures on the high seas survives fourth amendment scrutiny[ALL MODELS]
  • holding that inmates fourth amendment protection from unreasonable strip searches survives hudson
@@ 2.3. Zero-Shot Empirical Results
First, we measure the diversity of the model responses. When two models respond with the same answer to a given question, we code that as a match. The match rate measures how often two models give the same answer. In Figure 4, we show the match rate between the models. For example, the GPT-4o and NovaMicro models produced the same answer in 70.0% of cases. All model pairs produced the same answer more than 50% of the time.
@@
We measure the number of unique choices to a given question. Table 1 shows the number of unique choices for each question. In 34% of cases, all models tested chose the same answer - here, the models identified the correct choice more than 90% of the time. In other cases, the models split on the correct answer. Accuracy for questions with two or more different answers was much lower.
@@
We compute the macro F1 score which measures both precision (the fraction of correct predictions among all predictions of a given class) and recall (the fraction of actual instances of a class that were correctly identified). The macro F1 score averages the F1 score across all five answer choices to account for imbalance in how often each appears. Higher values are better, with 1 representing perfect performance. We include the best performing reported results from three papers examining the CaseHOLD dataset (Zheng et al., 2021; Chalkidis et al., 2022; Niklaus et al., 2024).
@@
  • As Figure 5 shows, performance scales with model size. Among the Llama, Nova and GPT-4o model families, performance on this task improves with parameter size. This aligns with the scaling laws seen elsewhere in the literature ((Brown et al., 2020; Kaplan et al., 2020)), and suggests that improvements to general purpose LLMs may improve performance on specific legal tasks.
  • Several models match or surpass the performance of custom-built legal models. GPT-4o performance of 0.744 surpasses the best results in (Zheng et al., 2021), 5 of the 7 results in (Chalkidis et al., 2022) (ranging from 0.708 to 0.754) and 3 of the 5 models presented in (Niklaus et al., 2024) (ranging from 0.717 to 0.770). These models do this without “fine-tuning” or domain adaptation. This means that general purpose LLMs may now take on tasks that historically required expensive and sophisticated fine-tuning.
@@ 3. Have LLMs memorized the test set?
To do so, we employ a two-step prompting scheme. In the first step, we use Prompt 6 with the GPT4o-mini model to anonymize the citation in question, replacing case names and citations with similar but artificial citations. The resulting text is semantically similar to the original, but modifications ensure that the text is novel and that the citations are meaningless. An example of such an original and modified citation prompt is presented in Figure 7. Note that all names and citations have been replaced with plausible alternatives, while the underlying facts and structure is unchanged.